### v1

# GraphCogent: Overcoming LLMs’ Working Memory Constraints via Multi-Agent Collaboration in Complex Graph Understanding

Abstract
Large language models (LLMs) show promising performance on small-scale graph reasoning tasks but fail when handling real-world graphs with complex queries. This phenomenon stems from LLMs’ inability to effectively process complex graph topology and perform multi-step reasoning simultaneously. To address these limitations, we propose GraphCogent, a collaborative agent framework inspired by human Working Memory Model that decomposes graph reasoning into specialized cognitive processes: sense, buffer, and execute. The framework consists of three modules: Sensory Module standardizes diverse graph text representations via subgraph sampling, Buffer Module integrates and indexes graph data across multiple formats, and Execution Module combines tool calling and model generation for efficient reasoning. We also introduce Graph4real, a comprehensive benchmark contains with four domains of real-world graphs (Web, Social, Transportation, and Citation) to evaluate LLMs’ graph reasoning capabilities. Our Graph4real covers 21 different graph reasoning tasks, categorized into three types (Structural Querying, Algorithmic Reasoning, and Predictive Modeling tasks), with graph scales that are 10 times larger than existing benchmarks. Experiments show that Llama3.1-8B based GraphCogent achieves a 50% improvement over massive-scale LLMs like DeepSeek-R1 (671B). Compared to state-of-the-art agent-based baseline, our framework outperforms by 20% in accuracy while reducing token usage by 80% for in-toolset tasks and 30% for out-toolset tasks. Code will be available after review.
1 Introduction
Large language models (LLMs) OpenAI (2023) have demonstrated remarkable cognitive capabilities in natural language processing. However, when querying the shortest path between two places in a 1000-node transportation graph, state-of-the-art LLMs like DeepSeek-R1 DeepSeek-AI (2025), GPT-o3 OpenAI (2023), and Gemini-2.5 pro Group (2023) make wrong turns in 9 out of 10 cases. This failure reveals that current LLMs remain limited in handling large real-world graph reasoning tasks.
Researchers have explored various methods to address the constraints of LLMs’ inability to perform graph reasoning. Text-based methods Fatemi et al. (2024); Luo et al. (2024); Chen et al. (2024) leverage Chain-of-Thought (CoT) Wei et al. (2022) reasoning but struggle with the dynamic programming requirements of graph algorithms (e.g., Bellman-Ford Bellman (1958)), resulting in cascading errors in complex tasks. Tool-based methods Zhang (2023); Wang et al. (2025b) rely on external tools for computation, requiring predefined tools and rigid input formats (e.g., preprocessed files). These limitations hinder their ability to process diverse graph text representations, such as adjacency lists, symbolic notations, and linguistic descriptions Fatemi et al. (2024), thereby reducing their adaptability to real-world scenarios. Agent-based methods Li et al. (2024) employ multi-agent collaboration, often decomposing graph tasks into sequential stages via naive instruction prompts. However, these approaches are fundamentally limited by the complexity of code construction and the inherent memory constraints of LLMs. As graph scales increase and tasks involve real-world graph problems, agent-based methods perform poorly when simultaneously handling graph data comprehension and task execution.
This limitation stems from LLMs’ working memory constraints that restrict simultaneous processing of complex graph topology and multi-step reasoning. A common mitigation approach involves partitioning graphs into smaller units for sequential processing Jiang et al. (2023), mirroring how humans decompose complex problems into manageable information chunks Miller (1956). Classic human working memory capacity assessments N-back test Kirchner (2006); Klatzky et al. (2008); Amon and Bertenthal (2018) (exampled in Fig. 1, left) reveals a rapid cognitive decline when humans are required to retain more than three items. Inspired by the N-back test, we design an analogous LLMs’ Graph N-back test. As exampled in the right of the Fig. 1, by the 3rd processing turn, Llama3.1-8B Touvron et al. (2023) exhibit significant topology forgetting, detailed quantitative result is analyzed in Section 3.1. This demonstrates that LLM graph reasoning similarly suffers from working memory constraints, mirroring human cognitive limitations.
Figure 1: Graph N-back Query Task: A graph is split into 50-edge subsets $E_{t}$. At turn $t+N$, the LLM verifies edge existence in $E_{t}$. Experimental results are in Section 3.1.
Human brain addresses these constraints through a specialized working memory model for sensory processing, information buffering, and task execution Baddeley et al. (2011). Facing a complex task, humans rely on an external sensory system to perceive stimuli, an episodic buffer to integrate and store information, and a central executive to coordinate cognition and processing. This absent architecture in monolithic LLMs reveals their key graph reasoning bottlenecks: (1) Diverse Text Representations: The sensory limitation in processing diverse graph text representations leads to inconsistent comprehension, especially for mainstream open-source models (e.g., Llama Touvron et al. (2023), Qwen Yang et al. (2024), and GLM GLM (2024)), thus impairing reasoning accuracy; (2) Overload Graph Scale: Lacking a buffer mechanism, LLMs’ limited input context windows hinders long-range dependency capture, causing global information loss and cascading reasoning errors; (3) Code Execution Fragility: The executive dysfunction results in unreliable code generation and inefficient algorithmic implementation.
Inspired by the cognitive architecture of human working memory system, we propose GraphCogent, an innovative Graph Collaboration Agentic framework designed to overcome LLMs’ working memory constraints in graph reasoning. Our framework consists of three modules: Sensory Module employs a sensory agent to sample subgraph and transform unstructured graph text representations into standardized adjacency list. Buffer Module integrates graph data from the Sensory Module and establishes data indices for diverse data formats (including Numpy, PyG, and NetworkX) based on Buffer Agent. Execution Module synergistically combines tool calling and model generation. Naive graph reasoning tasks are handled by a Reasoning Agent through a pre-built common toolset, while complex tasks leverage a Model Generation Agent that produces modular, task-specific components, avoiding error-prone full code generation by building upon the Buffer’s preprocessed data.
To evaluate LLM’s generalization capability in graph reasoning, we construct a more challenging Graph4real benchmark dataset, which collects graph data from four real-world domains: Web McAuley and Leskovec (2012), Social McAuley and Leskovec (2012), Transportation Shuman et al. (2013), and Citation Craven et al. (1998). Our Graph4real covers 21 different graph reasoning tasks, categorized into three types (Structural Querying, Algorithmic Reasoning, and Predictive Modeling tasks), with graph scales that are 10 times larger than existing benchmarks (e.g., NLGraph Wang et al. (2023), GraphWiz Chen et al. (2024)). Experimental results demonstrate that our method achieves state-of-the-art performance with an average accuracy of 98.5% based on the open-source Llama3.1-8B, representing a 20% improvement over other agent-based approaches. Cross-dataset validation demonstrates robust generalization, maintaining over 90% accuracy across diverse public benchmarks. The combining tool calling and model-generation strategy delivers dual optimization: it reduces token consumption by 80% for tasks covered by our toolset while still achieving 30% token savings for out-toolset tasks compared to the conventional agent-based method.
Our contributions are summarized as follows: (1) We propose GraphCogent, a novel agent collaboration framework addressing LLMs’ working memory constraints through its sensory-buffer-execution architecture; (2) We develop a combining reasoning strategy that integrates tool calling and model generation to handle diverse graph tasks while optimizing computational efficiency; (3) We construct Graph4real, a comprehensive benchmark featuring real-world graphs that covers 21 different graph reasoning tasks, categorized into three types (Structural Querying, Algorithmic Reasoning, and Predictive Modeling tasks), with graph scales that are 10 times larger than existing benchmarks.
2 Graph4real Construction
Current LLMs’ graph reasoning benchmarks face three key limitations: (1) Limited scale, primarily using randomly generated graphs that lack real-world topological characteristics; (2) Simple textual representations, over-relying on preprocessed files or adjacency lists while ignoring domain-specific semantic descriptions; (3) Artificial task formulations, where queries directly specify algorithms with only a single sentence rather than mirroring real-world reasoning scenarios requiring intent interpretation.
Real-World Graph Scaling. To establish a comprehensive benchmark for evaluating LLMs on real-world graph reasoning tasks, we curate a diverse set of graphs from four domains: Web (Google Web Graph McAuley and Leskovec (2012)), Social (SNAP Social Circles McAuley and Leskovec (2012)), Transportation (PeMS Shuman et al. (2013)), and Citation graphs (Cora Craven et al. (1998)). Unlike prior benchmarks Wang et al. (2023); Luo et al. (2024) that rely on randomly generated graphs, our dataset is constructed from real-world sources to ensure realistic topological properties and domain-specific characteristics. We employ a biased random walk sampling strategy to construct graphs at three scales (40, 100, and 1000 nodes), ensuring both effective evaluation of LLM-based methods through smaller-scale graphs and construct large-scale graphs to challenge existing methods.
Various Graph Text Representation. We formalize three graph text representation formats to evaluate LLMs’ graph comprehension and reasoning capabilities: (1) Adjacency list ([0,1],[0,2],…); (2) Symbolic notation (0→1, 2→3); (3) Linguistic descriptions using domain-specific predicates (Linked/Followed/Connected/Cited). These representations preserve various semantics, enabling assessment of LLMs’ ability to process both formal graph formats and domain-specific descriptions.
Intent-Driven Task Design. A total number of 21 tasks are designed and categorized into three classes: (1) Graph Structural Querying Tasks: Focus on fundamental graph properties, such as edge existence or node count, testing basic graph structural comprehension. (2) Graph Algorithmic Reasoning Tasks: Require reasoning based on classical algorithms, such as shortest path and maximum flow, evaluating algorithmic proficiency. (3) Graph Predictive Modeling Tasks: Involve neural network-based predictions, such as node classification and link prediction, assessing predictive modeling capabilities.
For task generation, we collect five real-world scenarios for each of the four domains (e.g., travel planning and logistics optimization for transportation) and craft 20 prompt templates to construct contextually grounded questions. Using DeepSeek-R1 DeepSeek-AI (2025), we generate graph reasoning tasks based on templates with specific scenarios, ensuring both diversity and relevance. To validate task quality, we employ a dual evaluation approach: DeepSeek-R1 for automated assessment and human for manual review. This process produces a dataset with 4200 questions distributed across the 21 tasks. The detailed task definitions, dataset statistics, and prompt templates are provided in Appendix F.
Figure 2: Overview of GraphCogent. Sensory Module (left) standardizes various graph text representations through subgraph sampling and conversion; Buffer Module (center) establishes cross-format data (e.g., NetworkX) integrating and indexing transformations; Execution Module (right) enables two reasoning modes: Reasoning Agent is employed for task discrimination and implements tool calling for in-toolset tasks, Model Agent handles out-toolset tasks based on model generation.
3 GraphCogent
We propose GraphCogent, an agent-based framework inspired by the human working memory system to address the challenges of real-world graph reasoning tasks. Our framework consists of three specialized modules: Sensory Module (samples subgraphs and transform unstructured graph text representations into standardized adjacency list), Buffer Module (integrates graph data from Sensory Module and establishes data indices for diverse data formats), and Execution Module (combines tool calling and model generation to comprehensively address graph reasoning tasks). The overall pipeline of the GraphCogent is illustrated in Fig. 2.
3.1 Sensory Module
Real-world graphs present two fundamental characteristics: (1) Large Graph Scale, benchmarks like NLGraph Wang et al. (2023) use small graphs (less than 40 nodes), while real-world graph (e.g., Cora Craven et al. (1998)) contains thousands of nodes and edges; (2) Varied Text Representation, different domains use distinct descriptions ("connected" in transportation or "followed" in social), and same text representation may appear in varied formats (adjacency lists, symbolic notations, linguistic descriptions).
Figure 3: Graph N-back Query. Accuracy (right y-axis) measures memory retention across dialogues, computed as sum of True edges (correctly identified existing edges) and False edges (correctly rejected non-existent edges) (left y-axis).
Figure 4: Graph Efficiency Coefficient (right y-axis) across sampling granularity. Edit shows graph Edit distance (left y-axis), and Cost (left y-axis, scaled ×100) counts total LLM interactions. Lower GEC indicates better performance.
We propose Graph N-back Query, a graph memory evaluation for LLM inspired by the human working memory capacity assessment through N-back tasks Kirchner (2006) (Detailed design of the Graph N-back Query and the working memory inspiration behind our framework are provided in Appendix A): Using a 100-node (sampled from PeMS datasets Shuman et al. (2013)) partitioned into 50-edge subsets $E_{t}$, we test memory retention across three text representations by querying edge existence at turn $t+N$. Experiments with Llama3.1-8B in Fig. 4 reveal significant information forgetting across dialogues, demonstrating that LLMs’ inherent memory mechanisms fail to reliably handle large-scale graph reasoning. This highlights that LLMs’ inherent context window limitations constrains their ability to maintain global graph information, rendering conventional step-by-step graph input approaches ineffective due to these working memory constraints.
The Sensory Module, which is proposed to address the dual challenges of graph scale and text representations, consists of two key components: Sensory Agent parses raw graph data via subgraph sampling and transform diverse text representations inputs into standardized adjacency lists; Graph Verifier serves as a supervisory mechanism for the Sensory Agent’s transformation process, detecting both format deviations and misaligned conversions.
Sensory Agent: Experiments in Fig. 4 demonstrate that the granularity of sampling directly impacts LLMs’ transformation performance. To achieve accurate text representation transformation, we define the Graph Efficiency Coefficient (GEC) to balance transformation accuracy (measured by Edit distance) against computational Cost (LLM token consumption):
$\text{GEC}=\text{Edit\_distance}(G,G^{\prime})\times\frac{\sum_{i=1}^{N}(T_{\text{input}}^{i}+T_{\text{output}}^{i})}{T_{\text{max}}},$ (1)
where $\text{Edit\_distance}(G,G^{\prime})$ measures the structural deviation between the original graph $G$ and the sampled graph $G^{\prime}$; $\sum_{i=1}^{N}(T_{\text{input}}^{i}+T_{\text{output}}^{i})$ quantifies the total token cost of LLM interactions, normalized by $T_{\text{max}}$ (maximum allowed tokens per LLM interaction). The algorithm for calculating Edit_distance and the rationale for selecting token count as the Cost metric are detailed in Appendix E.
We prepare 100 graphs (300 edges each) to test four subgraph granularities (25-100 edges). Using Llama3.1-8B, we measure Edit distance and Cost to compute GEC. Results in Fig. 4 show that coarser granularity exceeds LLM’s transformation capacity, increasing Edit distance, while finer granularity raises Cost through frequent LLM interactions. The GEC for graph reasoning tasks varies across models due to differences training or parameter scales, and for Llama3.1-8B, it achieves optimal GEC at approximately 50 edges.
Guided by GEC, the Sensory Agent decomposes large graphs into optimally-sized subgraphs to achieve accurate graph transformation. We employ heuristic prompts (prompt details in Appendix I) to guide the agent outputting standardized adjacency lists regardless of input format (symbolic notations, linguistic descriptions, etc.). To handle both weighted and unweighted graph adjacency lists, we implement dual regular expression templates to automatically match these two data formats. These templates precisely extract edge information from the parsed adjacency lists, with the extracted data subsequently validated by the Graph Verifier to ensure structural and semantic accuracy.
Graph Verifier: While GEC optimizes sampling granularity, LLM output uncertainty (shown by non-zero Edit distance in Fig. 4) may cause parsing errors during graph transformation. Since using standard graph data for correction risks information leakage, verification is constrained to known sampling granularity and target formats. The Verifier checks quantity consistency between original and transformed edges, and format compliance with adjacency list standards through regex parsing. Failed transformations trigger Sensory Agent retries.
3.2 Buffer Module
Current agent frameworks for graph reasoning lack storage mechanisms, relying instead on pre-processed files or delegating preprocessing to code generation. As Fig. 4 shows, even simple adjacency list transformations exhibit growing Edit distance with scale increasing, revealing LLMs’ declining reliability in graph comprehension. Moreover, graph paradigms diverge fundamentally in storage: algorithm tasks use OOP structures (e.g., NetworkX), link prediction requires tensor formats (e.g., PyG), and traffic prediction needs sparse matrices (e.g., NumPy). This difference overloads LLMs’ working memory, thus resulting in inaccurate code generation, demanding both task-accurate implementations and cross-format transformation via reasoning.
Based on the human episodic buffer mechanism Gong et al. (2024), we construct the Buffer Module to serve as an intermediary that not only stores and integrates graph data but also establishes indices for different data formats. This module transforms raw graph text representations into multiple standardized formats: NetworkX for graph algorithms, NumPy for numerical operations, and PyG for tensor requirement tasks.
The Buffer Module begins by integrating adjacency lists from the Sensory Module to construct complete graph structures. To support diverse reasoning tasks, the raw graph data is transformed into multiple standardized data formats. For Structural Querying and Algorithmic Reasoning tasks, the data is preprocessed into NetworkX graph objects that preserve topological relationships and node attributes. For Predictive Modeling tasks, we implement data preprocessing including normalization and outlier treatment before encapsulating the graph as NumPy arrays or PyG tensors. To enable efficient access across these data format, we implement the Buffer Agent to build data index that tracks key characteristics including data dimensionality, organizational schema, and metadata description. These indices are then passed to the Execution Module, allowing it to rapidly retrieve the most suitable graph format for each specific reasoning task.
3.3 Execution Module
Execution Module tackles code construction complexity and inefficiency challenges through two components: Reasoning Agent and Model Agent. The Reasoning Agent performs task analysis and tool calling using pre-built common toolset. It first evaluates whether a task can be solved using existing tools, then directly executes the appropriate tool when the task falls within the toolset’s coverage. For out-toolset tasks, it activates the Model Agent to generate specialized models. These models operate directly on the Buffer Module’s preprocessed data structures, avoiding the reliability issues of full code generation while maintaining the system’s adaptability to complex requirements.
3.3.1 Reasoning Agent
The Reasoning Agent’s tool calling and task discrimination capabilities are strengthened through a two-stage instruction tuning combining Thinking-Enhanced Supervised Fine-Tuning Hu et al. (2022) (SFT) with Direct Preference Optimization Rafailov et al. (2023) (DPO). This two-stage approach achieves both precise tool selection for in-toolset tasks and reliable discrimination of out-toolset scenarios.
Phase 1: Thinking-Enhanced Supervised Fine-Tuning for Tool Selection. We pair graph reasoning problems Prob with toolset descriptions Toolset, forming composite queries Ques = <Prob, Toolset>. These queries are processed by DeepSeek-R1 to generate outputs Ans = <Think, Tool>, where Think captures the reasoning process to improve decision transparency and tool selection robustness, while Tool indicates the selected tool. We curate high-quality training data by retaining only instances where the tool selection is correct, ensuring SFT process optimizes the following objective:
$\mathcal{L}_{\text{SFT}}=-\sum_{i=1}^{N}\log P(\langle\textit{Think, Tool}\rangle_{i}|\langle\textit{Prob, Toolset}\rangle_{i};\theta),$ (2)
where $N$ is the number of training examples, and $\theta$ represents the LLMs’ parameters.This phase enables the agent to develop active thinking and tool selection capabilities.
Figure 5: An overview of Reasoning Agent and Model Agent training process.
While Thinking-Enhanced SFT provides basic tool selection capabilities, Appendix C.3 reveals a critical limitation: the agent tends to force-fit out-toolset tasks into existing tools despite its active thinking abilities. This observation motivates our second-phase alignment approach.
Phase 2: Direct Preference Optimization Driven Coverage Recognition. To refine the agent’s tool coverage discrimination ability, we employ DPO to align its reasoning preferences. For each problem instance Ques, we generate multiple candidate responses from the SFT-tuned agent. These responses are then classified into two categories: Preferred responses ($R_{w}$): Correctly identify the out-toolset tasks with valid reasoning; Less preferred responses ($R_{l}$): Incorrectly associate tasks with tools. The DPO objective sharpens the agent’s decision boundary by optimizing:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(\text{Ques},R_{w},R_{l})}\left[\log\sigma\left(\beta\left(\log\frac{P_{\theta}(R_{w}|\text{Ques})}{P_{\text{SFT}}(R_{w}|\text{Ques})}-\log\frac{P_{\theta}(R_{l}|\text{Ques})}{P_{\text{SFT}}(R_{l}|\text{Ques})}\right)\right)\right],$ (3)
where $\beta$ controls deviation from the SFT-tuned model. This penalizes overconfidence in-toolset coverage and rewards accurate out-toolset detection. For each task, the Reasoning Agent first determines if it can be solved using the available toolset. If solvable, it constructs the tool calling using relevant Buffer Module data formats and executes the computation. Otherwise, it invokes the Model Agent for out-toolset processing.
3.3.2 Model Agent
The Model Agent handles out-toolset tasks by generating task-specific models. This maintains framework’s reliability through direct operation on the Buffer Module’s preprocessed data structures, avoiding full code generation issues. To develop the model generation capability, we construct a dataset consists of classical graph algorithms (e.g., PageRank) and neural models (e.g., GCN, GAT) from GitHub. Using LLMs (DeepSeek-v3 DeepSeek-AI (2025), GPT-4o OpenAI (2023), Claude-3.7 Anthropic (2024)), we generate executable graph reasoning codes. For each successful execution, we extract the core reasoning model component, creating training pairs (Ques, Model) to fine-tune Llama for model generation.
$\mathcal{L}_{\text{SFT}}=-\sum_{i=1}^{N}\log P(\textit{Model}_{i}|\textit{Ques}_{i};\theta),$ (4)
where $N$ represents validated examples and $\theta$ denotes LLMs’ parameters. This phase enables agent to generate task-specific models based on tasks while ensuring compatibility with preprocessed data. For out-toolset tasks from Reasoning Agent, Model Agent generates task-specific models based-on task descriptions and Buffer Module data formats. The framework integrates preprocessed data with generated models into executable code and produces final results.
3.3.3 Training Settings
We use Llama3.1-8B as the backbone for all agents, fine-tuned with LoRA Hu et al. (2022). Both Reasoning Agent and Model Agent are trained on Graph4real dataset. The training set employs 4 of 20 prompt templates, with their corresponding tasks excluded from the test set to ensure fair evaluation.
4 Experiments
4.1 Experimental Settings
Dataset. We evaluate on Graph4real, a real-world benchmark with graphs from 40 to 1,000 nodes spanning transportation, social, web, and citation domains. The dataset includes both in-toolset and out-toolset tasks, with 500 test instances per task and scale (across domains). To ensure fair comparisons, we use standard adjacency lists for graph text representation for main experiments in Table 1. Additional experiments on public datasets are provided in Appendix D.
Baseline Methods. We compare three types of approaches: (1) Text-based methods including GPT-4o OpenAI (2023), Claude-3.7 Anthropic (2024), and DeepSeek-R1 DeepSeek-AI (2025) with 2-shot CoT prompting. (2) Tool-based methods including Graph-Toolformer Zhang (2023), GraphTool-Instruction Wang et al. (2025b) on Llama3.1-8B Touvron et al. (2023) and GPT-4o with Function Calling OpenAI (2023); (3) Agent-based methods featuring the state-of-the-art GraphTeam Li et al. (2024) framework with GPT-4o-mini OpenAI (2023). All LLMs’ versions and settings are presented in Table 6.
Evaluation Metrics. We use accuracy for all tasks except traffic flow prediction (measured by MAE). Detailed metric computations are in Appendix D.
Table 1: Performances of GraphCogent and other baselines on Graph4real in-toolset tasks.
Method Model Edge Existence Edge Count Triangle Count Shortest Path Path Existence Cycle Detection Node Existence Node Count Degree Count Average
Small-Scale (40 Nodes)
Text GPT-4o OpenAI (2023) 100 22.2 20.6 25.8 77.2 78.2 100 84.2 29.8 59.8
Claude-3.7 Anthropic (2024) 99.2 22.4 21.0 21.2 79.8 81.8 100 80.2 32.2 59.8
DeepSeek-R1 DeepSeek-AI (2025) 100 32.6 26.6 30.8 86.6 86.8 100 90.4 31.0 65.0
Tool Graph-TF Zhang (2023) 48.0 75.6 71.0 63.8 44.2 66.8 50.2 77.2 50.0 60.8
GraphTool-Ins Wang et al. (2025b) 94.2 98.0 97.8 98.2 96.8 99.0 100 100 97.8 98.0
GPT4o-FC OpenAI (2023) 95.0 99.2 98.4 99.0 99.2 100 100 100 99.0 98.9
Agent GraphTeam Li et al. (2024) 86.6 84.8 82.8 80.2 97.6 80.0 96.8 70.8 97.6 86.4
GraphCogent 99.4 98.2 99.0 98.6 97.8 100 99.2 100 100 99.1
Middle-Scale (100 Nodes)
Text GPT-4o OpenAI (2023) 90.4 6.8 5.4 7.2 54.0 67.6 99.8 58.6 6.8 44.1
Claude-3.7 Anthropic (2024) 93.0 6.2 4.8 6.8 55.8 73.0 99.2 52.2 7.2 44.2
DeepSeek-R1 DeepSeek-AI (2025) 98.0 7.4 6.2 8.0 60.6 76.6 100 60.4 7.4 47.2
Tool Graph-TF Zhang (2023) 43.2 68.0 63.8 57.4 39.8 60.0 45.2 69.4 45.0 54.6
GraphTool-Ins Wang et al. (2025b) 84.8 89.0 88.0 88.4 89.0 89.0 90.6 92.0 88.0 88.8
GPT4o-FC OpenAI (2023) 85.4 89.2 88.6 89.0 89.2 89.4 97.0 94.2 89.0 90.1
Agent GraphTeam Li et al. (2024) 91.6 80.2 73.0 65.8 93.2 69.6 89.6 65.4 95.2 80.4
GraphCogent 98.2 97.8 98.8 98.2 97.4 99.0 99.4 97.6 99.8 98.5
Large-Scale (1000 Nodes)
Agent GraphCogent 97.0 97.4 98.6 98.0 97.2 97.0 96.8 97.2 99.6 97.6
Note: Graph-TF, GraphTool-Ins ,and GPT4o-FC are abbreviations for Graph-Toolformer, GraphTool-Instruction, and GPT4o-Function Calling.
4.2 Main Result
Figure 6: Code execution rate on out-toolset tasks.
The performance results are reported in Table 1. Our method significantly outperforms all baseline methods across toolset-covered tasks. As for text-based methods, they demonstrate advantages on small-scale simple Structural Querying tasks (e.g., DeepSeek-R1 achieves 100% on edge existence). However, these methods show significant performance degradation when dealing with problems requiring multi-step reasoning (e.g., only 30.8% accuracy on shortest path), and this issue becomes more pronounced as the scale increases. Tool-based approaches such as GPT4o-FC demonstrates robust performance across all tasks on small-scale graphs by leveraging GPT4o’s powerful reasoning capabilities. However, they suffer 10% performance degradation when scaling from 40 to 100 nodes. The agent-based GraphTeam demonstrates limitations in code generation across multiple tasks. The complex data preprocessing and reasoning requirements overload the model’s working memory, resulting in lower performance. In contrast, our method maintains state-of-the-art performance across all scales with less than 1% variance. Notably, while no baseline could handle large-scale graphs, we have only included the results of our approach.
Table 2: Performance evaluation on out-toolset tasks.
Transportation Social Web Citation
Method Max Diameter Traffic Maxcore Connect. Link Common PageRank Link Reference Cluster. Node
Flow Calcu. Pred. Calcu. Compo. Pred. Neighbor Calcu. Pred. Match Coeff. Class.
GraphTeam 72.4 61.6 96.7 69.8 72.0 76.4 51.2 76.8 71.6 78.8 76.4 72.0
GraphCogent 89.6 92.4 35.1 91.6 87.6 85.2 88.8 92.6 86.4 90.0 85.2 89.6
Note: Except for the traffic prediction (which values represent MAE), all other values represent Accuracy.
Table 3: Reasoning token comparison between GraphTeam and GraphCogent.
In-Toolset Tasks Out-Toolset Tasks
Method Shortest Path Cycle Detection Max Flow Link Prediction
Input Output Accuracy Input Output Accuracy Input Output Accuracy Input Output Accuracy
GraphTeam 4248 3285 65.8 4207 2901 69.6 4122 3023 72.4 4423 3820 76.4
GraphCogent 2794 615 98.2 2533 595 99.0 3009 1044 89.6 3128 1244 90.2
Note: Input and Output are abbreviation for average input tokens and average output tokens respectively.
We further evaluate GraphCogent’s performance on out-toolset tasks against GraphTeam in Table 2 on middle-scale graphs. Focusing on agent capabilities for handling uncovered tasks, we omit tool-based methods from comparison as they inherently lack these functionalities. Our method demonstrates superior adaptability, achieving over 18% improvements over GraphTeam across four domains.
Experiments on effectiveness evaluation of our approach demonstrate the superiority in two key aspects. First, in terms of code executability shown in Fig. 6, our method achieves higher one-time success rates compared to GraphTeam’s three-retry mechanism, validating the robustness of our Execution Module. Second, our token consumption analysis in Table 3 shows consistent improvements across all tasks, with significantly reduced input token consumption and substantial output token savings (approximately 80% for in-toolset tasks and 30% for out-toolset tasks) while maintaining high accuracy.
Table 4: Reasoning time comparison between GraphCogent and other baselines.
Methods Shortest Path Cycle Detection Triangle Count Path Existence Link Prediction
Time Acc. Time Acc. Time Acc. Time Acc. Time Acc.
GPT4o-Function Calling 12.2s 89.0 10.7s 89.4 11.3s 88.6 12.9s 89.2 / /
Graph-Toolformer 18.1s 57.4 16.6s 60.0 18.9s 63.8 17.3s 39.8 / /
GraphTool-Instruction 19.8s 88.4 18.3s 89.0 20.1s 88.0 19.2s 89.0 / /
GraphTeam 58.1s 65.8 56.6s 69.6 57.3s 73.0 58.9s 93.2 59.5s 76.4
GraphCogent 22.8s 98.2 23.3s 99.0 20.1s 98.8 19.2s 97.4 24.0s 85.2
We compare Tool-based and Agent-based methods on reasoning time in Table 4. While GPT4o-FC’s API implementation demonstrates significant time advantages over local open-source LLMs, our local method GraphCogent achieves competitive speed while maintaining superior accuracy. The additional reasoning time in our framework mainly stems from multi-round sampling in the Sensory module. In contrast, although GraphTeam (based on GPT4o-mini) benefits from faster API inference, its performance is compromised by working memory overload in complex tasks and the overhead of its three-retry mechanism, making its direct code generation approach both slower and less accurate.
Table 5: Accuracy on larger graph
Scale Shortest Cycle Triangle Path Path Detect. Count Exist. 2000 97.0 97.0 98.6 97.0 5000 98.2 97.0 98.8 98.0 10000 97.4 96.8 99.0 97.0
The scalability evaluation in Table 5 on larger graphs further demonstrates the effectiveness of our framework, where it maintains stable performance across various tasks. Through ablation studies in Appendix C and adaptability analysis in Appendix D, we examine the contributions of each component and demonstrate the adaptability of our method to based on different LLMs.
5 Related work
LLMs for Graph Reasoning Tasks. Existing graph reasoning methods adopt three paradigms: Text-based Fatemi et al. (2024); Luo et al. (2024); Chen et al. (2024); Guo et al. (2023); Wang et al. (2024) approaches leverage chain-of-thought prompting to decompose graph algorithms stepwise, but suffer error accumulation in multi-step reasoning. Tool-based Zhang (2023); Wang et al. (2025b); OpenAI (2023) methods offload computations to external solvers (e.g., NetworkX), yet rigid input requirements limit adaptability to diverse text representations. Agent-based Li et al. (2024) frameworks attempt task decomposition through multi-agent collaboration, but face scalability bottlenecks when handling large graphs due to code generation complexity and memory overload.
Working Memory Constraints in LLMs. Like humans facing working memory constraints in complex tasks Jaeggi et al. (2008); O’Donohue et al. (2003), LLMs show similar capacity constraints. Cognitive studies show humans typically retain only few information units simultaneously, with performance decaying sharply in N-back tests (where N = 3) Kirchner (2006); Klatzky et al. (2008); Amon and Bertenthal (2018); Jaeggi et al. (2010); Wang et al. (2025a). LLMs mirror this pattern, exhibiting accuracy declines on 3-step dependencies Baddeley and Hitch (1974). This working memory overload manifests in graph reasoning as excessive graph scale and complex reasoning tasks, thus hindering effective processing of intricate graph structures and multi-step reasoning.
Graph Reasoning Benchmarks for LLMs. NLGraph Wang et al. (2023), as the pioneering benchmark, demonstrates LLMs’ competence on small-scale graphs (less than 40 nodes) with basic reasoning tasks. GraphWiz Chen et al. (2024) later expands to 100 nodes with more complex tasks. However, these and other benchmarks Luo et al. (2024); Wang et al. (2024); Fatemi et al. (2024); Wang et al. (2025b) remain constrained by fixed-size graphs within context windows and templated reasoning tasks, failing to match real-world graph reasoning demands where both scale and complexity exceed current test conditions.
6 Conclusion
In this work, we proposed GraphCogent, a novel agent framework that enhances LLMs’ graph reasoning capabilities by addressing their working memory constraints. Inspired by human cognitive architecture, GraphCogent integrates three key modules: Sensory Module (standardizing diverse graph text representations), Buffer Module (integrating and indexing graph data), and Execution Module (combining tool-based and model-based reasoning). To evaluate LLMs on real-world graph reasoning tasks, we introduce Graph4real, a benchmark featuring large-scale graphs for four real-world domains. Experiments show that GraphCogent achieves 20% higher than existing state-of-the-art agent-based method while reducing over 30% token usage.
Appendix A Working Memory Inspiration
Figure 7: The correspondence between the human working memory model and GraphCogent.
The design of GraphCogent is based on the human working memory model Baddeley et al. [2011]. We start with a key observation: while LLMs excel at natural language processing, they struggle with complex graph tasks like finding shortest paths in transportation networks. A common mitigation approach for LLMs to reason on large graph is partitioning graphs into smaller units for sequential processing Jiang et al. [2023], mirroring how humans decompose complex problems into manageable information chunks. However, classic human working memory capacity assessments N-back test Kirchner [2006], Klatzky et al. [2008], Amon and Bertenthal [2018] reveals a rapid cognitive decline when humans are required to retain more than three items.
In a typical Verbal N-back experiment Klatzky et al. [2008], participants are presented with a sequence of characters (e.g., H, B, D, A, E, D, …) and asked to determine whether the current character matches the one presented $N$ steps earlier. For instance, in a 3-back task, the correct response at position $t+3$ would depend on recalling the character at position $t$. Humans typically show significant accuracy drops when tracking information beyond three delayed turns ($N=3$). This indicates that the decomposition of complex tasks into manageable information chunks, which humans rely on, has inherent limits. Exceeding these working memory constraints leads to marked cognitive decline.
Inspired by the N-back test, we design an analogous Graph N-back test for LLMs to investigate whether they similarly encounter working memory constraints when performing graph reasoning tasks. We randomly sample 100 graphs (each containing 100 nodes with 300 edges) from PeMS dataset, partitioning each graph into 50-edge subsets $E_{t}$, with different graph text representations (i.e., adjacency list, symbolic notation, and linguistic description). At turn $t+N$, the model is queried about the existence of specific edges in $E_{t}$. For example, when $N=3$, the model must recall edges from three turns earlier to verify their presence or absence. Results reveal a parallel: LLMs like Llama3.1-8B demonstrates sharp accuracy declines at $N=3$ across all graph text representations, mirroring human cognitive decay patterns (Fig. 9). We further test diverse open-source LLMs (Qwen2.5-7B and GLM4-9B) in Fig. 9 at adjacency list, while their absolute performance varied due to differences in training methodologies and parameter counts, all exhibit consistent accuracy deterioration trends as $N$ increased. This alignment suggests that LLMs’ failures in graph tasks analogous to human working memory constraints.
Figure 8: Graph N-back Query on Llama3.1-8B at different text representations. Accuracy (right y-axis) measures memory retention across dialogues, computed as sum of True edges (correctly identified existing edges) and False edges (correctly rejected non-existent edges) (left y-axis).
Figure 9: Graph N-back Query on Llama3.1-8B, GLM4-9B, and Qwen2.5-7B at graphs represented in adjacency list. Accuracy (right y-axis) measures memory retention across dialogues, computed as sum of True edges and False edges (left y-axis).
The observed working memory constraints in both humans and LLMs reveal fundamental limitations in concurrent information maintenance and processing. Cognitive studies Baddeley and Hitch [1974] demonstrate that humans mitigate these constraints through specialized systems named working memory model. The working memory model consists of three components: Central executive for coordination and processing; Phonological loop and Visuospatial sketchpad for modality-specific processing (collectively termed External Sensory); Episodic buffer for integrating and temporarily storing information. This biological solution inspires our artificial system design: rather than seeking to expand absolute memory capacity, we emulate the human strategy of functional decomposition and specialized processing. Mapping this cognitive architecture into our framework, we have identified that existing LLMs’ graph reasoning methods whatever tool-based or agent-based fail due to following three key gaps: (1) Sensory limitations in processing diverse graph text representations (e.g., adjacency lists, symbolic notations, linguistic descriptions), leading to inconsistent comprehension; (2) Buffer deficiencies, where the absence of a storage mechanism to integrate and index graph data across formats results in information loss and cascading reasoning errors; (3) Executive dysfunction, manifesting as unreliable code generation and inefficient algorithmic implementation due to monolithic reasoning processes.
Based on these insights, we propose GraphCogent, which aims to address working memory constraints by emulating the architecture of human working memory models. Our three core modules directly correspond to the three fundamental components of human working memory model and effectively resolve the three key challenges. The experimental results demonstrate the efficacy of our approach, validating that this biologically inspired design successfully overcomes the inherent limitations of LLMs in graph reasoning tasks.
Appendix B Experimental Settings
Dataset. We evaluate our approach on Graph4real, a real-world benchmark designed for LLM-based graph reasoning tasks, with graph sizes ranging from 40 to 1,000 nodes and spanning four real-world senarios transportation, social, web, and citation. The dataset includes both toolset-covered tasks (including all Structural Querying tasks and four Algorithmic Reasoning tasks) and out-toolset tasks (including all Predictive Modeling tasks and eight Algorithmic Reasoning tasks). For each task and each scale, we generate 200 questions from each of the four domains and combine the questions with different graphs (sampled and generated based on random walk methods) to get 500 test instances. We calculate the average performance of various LLMs across these 500 test instances. To ensure fair comparisons, for the main experimental results in Table 1, we uniformly adopt standard adjacency lists to represent graph information, aligning with the default reasoning mode of most baseline methods. Additionally, we introduce public datasets: NLGraph Wang et al. [2023], Talk like Graph Fatemi et al. [2024], and GraphWiz Chen et al. [2024].
Baseline Methods. We compare three categories of approaches:
  • Text-based methods including current state-of-the-art LLMs (GPT-4o OpenAI [2023], Claude-3.7 Anthropic [2024], and DeepSeek-R1 DeepSeek-AI [2025]) with 2-shot CoT prompting. Notably, since existing open-source graph reasoning methods like GraphWiz Chen et al. [2024] only demonstrate reasonable performance under their specific training paradigms, we exclude them from main comparisons.
  • Tool-based methods including reproduced Graph-Toolformer Zhang [2023] and GraphTool-Instruction Wang et al. [2025b] based on Llama3.1-8B Touvron et al. [2023] as the representative of open-source approaches. Additionally, we employ GPT-4o OpenAI [2023] with function calling as the closed-source baseline.
  • Agent-based methods featuring the state-of-the-art GraphTeam Li et al. [2024] framework with GPT-4o-mini OpenAI [2023], maintaining the same configuration as in the original work.
All corresponding versions of the LLMs are presented in Table 6.
Table 6: Summary of baseline methods and corresponding models.
LLM Type Method Base Model
Text-based Two-shot prompt Wei et al. [2022] GPT-4o-2024-08-06 OpenAI [2023]
Two-shot prompt Wei et al. [2022] Claude-3.7-sonnet-20250219 Anthropic [2024]
Two-shot prompt Wei et al. [2022] DeepSeek-R1 DeepSeek-AI [2025]
Tool-based Graph-Toolformer Zhang [2023] Llama3.1-8B-Instruct Touvron et al. [2023]
GraphTool-Instruction Wang et al. [2025b] Llama3.1-8B-Instruct Touvron et al. [2023]
Function Calling OpenAI [2023] GPT-4o-2024-08-06 OpenAI [2023]
Agent-based GraphTeam Li et al. [2024] GPT-4o-mini-2024-07-18 OpenAI [2023]
GraphCogent Llama3.1-8B-Instruct Touvron et al. [2023]
Evaluation Metrics. For the Graph4real dataset, we employ accuracy as the unified evaluation metric for all Structural Querying, Algorithmic Reasoning, and Predictive Modeling tasks (excluding traffic prediction), we use accuracy as the unified evaluation metric. For the traffic prediction task, we report its Mean Absolute Error (MAE). The detailed metric computation methods for each task category are provided in Appendix D. For public datasets: NLGraph Wang et al. [2023], Talk like Graph Fatemi et al. [2024], and GraphWiz Chen et al. [2024]. We directly compute the overall accuracy of each method across all tasks and present the results in Table 12.
Setting. All open-source LLM-based methods were trained on 8*NVIDIA 80G A800 GPUs and evaluated on 8*NVIDIA 24G 4090 GPUs. The LLMs’ primary parameters were configured with default values, maintaining a temperature of 0.7 and top_p of 1. Both LoRA and DPO training were implemented using the Llama-Factory framework. For LoRA training, we employed a learning rate of 1e-5 with a 0.1 warmup ratio, a batch size of 4, and a cosine learning rate scheduler for cyclical adjustments over 5 epochs. The DPO training utilized a batch size of 1, 4 gradient accumulation steps, a learning rate of 5.0e-6, and 5 training epochs with cosine learning rate scheduling and a 0.1 warmup ratio. Closed-source LLMs were accessed through the official APIs of OpenAI and Anthropic, with their parameters set to default values temperature with 0.7 and top_p with 1.
Appendix C Ablation Study
We conduct systematic ablation studies to evaluate the contributions of each module and their key components through the following research questions:
  • Sensory Module
    • RQ1: How does the Graph Verifier component enhance transformation reliability?
  • Buffer Module
    • RQ2: How does the Buffer Mechanism enhance the framework’s reasoning capability?
  • Execution Module
    • RQ3: Does Thinking-Enhanced SFT improve Reasoning Agent’s tool selection?
    • RQ4: Why employ DPO for coverage discrimination in Reasoning Agent?
    • RQ5: How does Model-Enhanced improve Model Agent’s generation capability?
C.1 Ablation Studies on Sensory Module
RQ1: How does the Graph Verifier component enhance transformation reliability? As shown in Fig. 4) even based on GEC optimization, the inherent uncertainty in LLM outputs can still introduce parsing inaccuracies during graph transformation. Table 7 further demonstrates the critical role of the Graph Verifier in ensuring reliable graph transformation and standardization.
We randomly sample 100 graphs (each containing 100 nodes with 300 edges) from PeMS dataset, partitioning each graph into 50-edges subsets with different graph text representations (i.e., adjacency list, symbolic notation, and linguistic description). We measure the total edit distance during the transformation phase. The results show that introducing the Graph Verifier reduces Edit distance by approximately 90%, while maintaining operational efficiency with fewer than 10 verification triggers per transformation. The results show that introducing the Graph Verifier reduces structural errors by approximately 90%, while maintaining operational efficiency with fewer than 10 verification triggers. This retry frequency (occurring in approximately 600 LLM calls) remains acceptable while significantly reducing edit distance.
Table 7: Impact of Graph Verifier.
Representation Type Edit Distance (Before) Edit Distance (After) Verifier Triggers
Adjacency List 44 5 6
Symbolic Notation 57 6 9
Linguistic Description 66 5 9
C.2 Ablation Studies on Buffer Module
RQ2: How does the Buffer Mechanism enhance the framework’s reasoning capability? We select four tasks within Graph4real on middle-scale: Shortest Path, Cycle Detection, Max Flow, and Link Prediction. The experimental results in Table 8 show that removing the Buffer Module causes substantial performance decreases across all tasks. This finding evidences the Buffer Module’s critical role in addressing LLMs’ working memory constraints for graph reasoning. Without the Buffer Module, when the graph scale exceeds the capacity of LLMs’ working memory, the LLMs exhibit substantial information loss regarding graph topology.
Table 8: Performance impact of Buffer Module removal.
Task Accuracy (%)
Variants Shortest Path Cycle Detection Max Flow Link Prediction
GraphCogent 98.2 99.0 89.6 85.2
GraphCogent w/o Buffer 68.4 69.2 53.8 54.2
C.3 Ablation Studies on Execution Module
RQ3: Does Thinking-Enhanced SFT improve Reasoning Agent’s tool selection? We select four tasks within Graph4real on middle-scale: Shortest Path, Cycle Detection, Max Flow, and Link Prediction. Three versions of the Reasoning Agent are evaluated: the original untuned agent (Llama3.1-8B), agent fine-tuned using the Naive method, and agent fine-tuned with our Thinking-Enhanced method. The Naive fine-tuning approach involves training Llama3.1-8B by directly mapping problem statements to their outputs, where each output consists of a brief analytical statement containing the tool selection result. As shown in Table 9, the Thinking-Enhanced SFT method shows consistent performance advantages across different tasks. For in-toolset tasks such as Shortest Path and Cycle Detection, Thinking-Enhanced SFT achieves a 5% performance improvement over the naive approach, as the explicit reasoning chains enable the model to better understand task intent and select appropriate tools. Through analysis of the reasoning process, we observe that compared to Naive SFT’s tendency to directly output corresponding tools (i.e., simply fitting tool results), the Thinking-Enhanced approach teaches the agent to comprehend task intentions and demonstrates reasonable reasoning paths, leading to correct tool selection. This advantage becomes even more significant for out-toolset tasks, where the Thinking-Enhanced method shows substantially better generalization capabilities than naive method.
Table 9: Impact of Thinking-Enhanced SFT on Reasoning Agent performance.
Task Accuracy (%)
Training Method Shortest Path Cycle Detection Max Flow Link Prediction
Llama3.1-8B (Base) 78.4 79.2 35.8 36.2
Llama3.1-8B w/ Naive SFT 89.3 90.2 43.8 44.2
Llama3.1-8B w/ Thinking SFT 98.2 99.0 70.2 65.0
However, the Thinking-Enhanced approach still has limitations. Through analyzing the reasoning results for tasks like Max Flow and Link Prediction, we identify a tendency of the LLM to forcibly associate tools. For instance, although no single tool can solve the Max Flow problem, unlike the Naive SFT-tuned LLM, which randomly generates a tool or calls a non-existent tool, the Thinking-Enhanced SFT-tuned LLM attempts to combine tools to derive a solution. This preference actually contradicts practical decision-making. In this task, we would prefer the agent to directly generate the task-specific model of Max Flow through Model Agent rather than repeatedly interact with tools of Reasoning Agent. This observation motivates our implementation of DPO alignment to refine the Reasoning Agent’s preferences.
RQ4: Why employ DPO for coverage discrimination in Reasoning Agent? Building upon the findings from C.3, the core objective of DPO alignment is to refine the Reasoning Agent’s tool selection preferences for improving toolset discrimination capabilities. Although tasks are not strictly binary (i.e., solvable either by tools or not at all), the LLM requires a decision-making mechanism to optimize resource efficiency. The goal is to minimize resource consumption, single tool calls can significantly reduce the LLM’s interaction tokens, whereas chained tool calls may not be as advantageous as direct model generation. The experimental results in Table 10 demonstrate that DPO not only preserves Reasoning Agent’s in-toolset selection performance but also substantially enhances out-toolset discrimination capability.
Table 10: Impact of DPO on Reasoning Agent.
Task Accuracy (%)
Training Method Shortest Path Cycle Detection Max Flow Link Prediction
Reasoning Agent w/o DPO 97.2 97.8 70.2 64.8
Reasoning Agent w/ DPO 98.2 99.0 89.7 85.2
RQ5: How does Model-Enhanced SFT improve Model Agent’s generation capability? We select four out-toolset tasks within Graph4real on middle-scale: Max Flow, Commoin Neighbors, Link Prediction, and Node Classification. Three versions of the Model Agent are evaluated: the original untuned agent (Llama3.1-8B), agent fine-tuned using Code-Enhanced SFT method, and agent fine-tuned using Model-Enhanced SFT method. Table 11 reports the accuracy and executable percentage. The improvements in both accuracy and executability metrics confirm the effectiveness of Model-Enhanced SFT. Compared to Code-Enhanced SFT, especially for mainstream open-source LLMs, it is more challenging to enable LLMs to master the reasoning capabilities for complex tasks through code training. The complexity of the tasks and the limitations of LLM capabilities hinder their generalization performance on complex tasks. In contrast, Model-Enhanced SFT effectively enhances accuracy and executability by allowing the model to focus more on generating core task-specific models.
Table 11: Impact of Model-Enhanced SFT on Model Agent’s performance.
Without SFT With Code SFT With Model SFT
Task Acc.(%) Execut.(%) Acc.(%) Execut.(%) Acc.(%) Execut.(%)
Max Flow 38.2 58.4 72.2 81.2 89.6 96.3
Common Neighbors 42.6 63.1 70.6 78.3 88.8 95.8
Link Prediction 39.8 61.6 58.6 65.2 85.2 88.4
Node Classification 41.2 59.8 60.8 64.7 89.6 89.9
Appendix D Public Dataset Performance and Framework Adaptability
D.1 Performance on Public Datasets
Table 12: Performance comparison across graph reasoning benchmarks.
Method Talk like a Graph GraphWiz NLGraph Average
GPT4o-FC 100 99.61 100 99.87
GraphTool-Instruction 98.20 98.02 99.70 98.64
GraphTeam 99.13 88.61 97.80 95.18
GraphCogent 99.30 90.02 99.70 96.34
We conduct a systematic comparison between tool-based and agent-based approaches using public benchmarks in Table 12. For fair evaluation, we use GraphTool-Instruction’s predefined toolset as the common toolset and report tool-based methods’ performance within in-toolset tasks. The results demonstrate that GraphCogent maintains strong performance across all three public datasets. Compared to the agent-based baseline GraphTeam, our approach consistently outperforms across all three datasets. For Tool-based methods, especially GPT4o-Function Calling, achieve nearly perfect accuracy on supported tasks. This high performance stems from GPT-4o’s strong reasoning capabilities. However, this tool-based approach cannot process queries requiring tools outside their predefined set, significantly limiting their flexibility.
D.2 Framework Adaptability across Different LLMs
Table 13: Framework adaptability across different LLMs.
Task Accuracy (%)
Base Model Shortest Path Cycle Detection Max Flow Link Prediction
Naive GLM4-9B 0.0 68.4 0.0 48.6
GraphCogent (GLM) 96.2 96.8 81.4 82.2
Naive Qwen2.5-7B 0.0 56.4 0.0 51.0
GraphCogent (Qwen) 92.2 90.4 91.2 87.2
Naive GPT-4o 7.2 67.6 10.2 62.4
GraphCogent (GPT-4o) 100 100 99.2 90.8
Naive DeepSeek-R1 8.0 76.7 9.6 60.4
GraphCogent (DeepSeek-R1) 100 100 99.2 91.2
Our experiments primarily focus on the Llama3.1-8B, but the framework’s modular design ensures compatibility with both open-source and closed-source LLMs, as confirmed by cross-LLM validation results. We select four tasks within Graph4real on middle-scale: Shortest Path, Cycle Detection, Max Flow, and Link Prediction. Table 13 reveals our framework’s adaptability. For open-source models like GLM4-9B-chat and Qwen2.5-7B-coder-Instruct, our framework significantly enhances their reasoning capabilities especially for complex tasks requiring multi-step graph operations, improving performance from near-zero accuracy in tasks such as Shortest Path and Max Flow to high accuracy. Second, based on the framework’s instruction without training, massive-scale LLMs like GPT-4o and DeepSeek-R1 achieve high performance with in-toolset tasks and out-toolset, demonstrating the universal applicability of our framework design.
Appendix E Edit distance and Cost for GEC
E.1 Edit distance
Real-world scenarios present graphs in multiple text representations including adjacency lists, symbolic notations, and linguistic descriptions. This variability makes rule-based preprocessing approaches such as regular expressions impractical for standardization. Mirroring how human external sensory systems process multimodal information, our Sensory Module transforms these varied graph inputs into standardized adjacency list representations.
The transformation process requires rigorous quality assessment due to the substantial topological information contained in real-world graphs and the cascading effects of transformation errors on downstream reasoning. We employ edge Edit distance as shown in algorithm 1, quantify the minimum number of edge additions and deletions required to transform one graph into another to evaluate difference between two graphs.
E.2 Cost for Graph Efficiency Coefficient
The selection of token count as the primary cost metric in GEC computation is motivated by three considerations. First, token-based pricing serves as the universal billing standard across LLM platforms (e.g., GPT, Claude, and DeepSeek series), making it the most feasible cost indicator for comparative studies. Second, while temporal cost was initially considered, our experiments with 100 graphs (300 edges each) across four sampling granularities revealed less than 5% variance in processing time. This temporal consistency emerges because the output lengths remain relatively stable for graph transformations. Crucially, the actual computational expense manifests in input token caused by LLMs’ input instruction (i.e., the instruction guides the LLM to transform graph information into a fixed adjacency list format). We therefore adopt normalized token counts as they precisely capture this trade-off: finer sampling granularity necessitates more frequent LLM interactions, each accumulating input instruction tokens, while coarser granularity reduces interaction frequency at the risk of topological distortion.
Algorithm 1 Graph Edit Distance Computation
1:
2: Source graph $G=(V,E)$
3: Target graph $H=(V,E^{\prime})$ with identical node set $V$
4:
5: Edit distance $d_{\text{edit}}(G,H)$ between graphs $G$ and $H$
6:Initialize empty edge sets:
[…]
9:Convert edge lists to sets:
[…]
12:Compute edge differences:
13:for each edge $e\in E_{H}$ do
14:  if $e\notin E_{G}$ then
15:   $\mathcal{A}\leftarrow\mathcal{A}\cup\{e\}$
16:  end if
17:end for
18:for each edge $e\in E_{G}$ do
19:  if $e\notin E_{H}$ then
20:   $\mathcal{R}\leftarrow\mathcal{R}\cup\{e\}$
21:  end if
22:end for
23:Calculate final distance:
[…]
25:return $d_{\text{edit}}(G,H)$
Appendix F Graph Reasoning Task Definition
We design a total number of 21 tasks and categorize them into three classes as following:
  • Graph Structural Querying Tasks: Focus on fundamental graph properties, such as edge existence or node count, testing basic graph structural comprehension.
  • Graph Algorithmic Reasoning Tasks: Require reasoning based on classical algorithms, such as shortest path computation and maximum flow analysis, evaluating algorithmic proficiency.
  • Graph Predictive Modeling Tasks: Involve neural network-based predictions, such as node classification and link prediction, assessing predictive modeling capabilities.
Table 14: Domains and corresponding scenarios.
Domain Scenarios
Social
Information diffusion analysis
Community detection and recommendation systems
Fraudulent account detection
Influence maximization algorithms
Social network dynamics analysis
Web
Web crawler efficiency optimization
Search engine ranking optimization
Web structural integrity diagnosis
Topical community discovery
DDoS attack mitigation
Transportation
Travel route planning
Logistics delivery optimization
Urban emergency response planning
Public transit network scheduling
Shared mobility platforms
Citation
Scholarly influence tracking
Interdisciplinary research identification
Seminalpaper discovery
Literature retrieval ranking optimization
Research frontier identification
For task generation, we collect five real-world scenarios as shown in Table 14 for each of the four domains (e.g., travel planning and logistics optimization for transportation) and craft 20 prompt templates (a template example is shown in Fig. 14) to construct contextually grounded questions. Using DeepSeek-R1 DeepSeek-AI [2025], we generate graph reasoning tasks based on templates with specific scenarios, ensuring both diversity and relevance. To validate task quality, we employ a dual evaluation approach: DeepSeek-R1 for automated assessment and human for manual review. This process produces a dataset with 4200 questions distributed across the 21 tasks.
F.1 Graph Structural Querying Tasks
The Graph Structural Querying tasks in Table 15 represent fundamental operations that are universally applicable across all four domains (Web, Social, Transportation, and Citation) in our benchmark. These tasks focus on direct lookup operations with constant time complexity, avoiding complex algorithmic requirements to measure basic graph comprehension abilities. Using designed task templates, we generate domain-specific versions of each Structural Querying task that maintain the same core operation while incorporating appropriate context from each domain. The five tasks include Edge Count, Node Count, Degree Count, Edge Existence, and Node Existence. All these tasks are classified as in-toolset, with each task having a corresponding pre-built tool implementation that performs direct lookup operations on the graph data structure.
Table 15: Details of Graph Structural Querying Tasks.
Task Description Tool Algorithm Time Complexity
Edge Count Count the total number of edges in a given graph. Direct Lookup $O(1)$
Node Count Count the total number of nodes in a given graph. Direct Lookup $O(1)$
Degree Count Count the number of edges connected to a specific node in a given graph. Direct Lookup $O(1)$
Edge Existence Determine whether a specific edge exists between two nodes in a given graph. Direct Lookup $O(1)$
Node Existence Determine whether a specific node exists in a given graph. Direct Lookup $O(1)$
F.2 Graph Algorithmic Reasoning Tasks
The in-toolset Graph Algorithmic Reasoning Tasks in Table 16 extend beyond basic graph comprehension to assess multi-step reasoning capabilities. These tasks employ classical graph algorithms with varying time complexities, from linear to cubic, to evaluate different aspects of structural understanding. Tasks contain Cycle Detection, Triangle Count, Path Existence, and Shortest Path. Mirroring Structural Querying tasks, these algorithmic challenges are implemented as pre-defined tools and adapted cross-domains template for each scenario.
Table 16: Details of in-toolset Graph Algorithmic Reasoning Tasks.
Task Description Tool Algorithm Time Complexity
Cycle Detection Determine whether there exists any cycle in a given graph. Depth-First Search $O(|V|+|E|)$
Triangle Count Count the total number of triangles in a given graph. Node Iterator $O(|V|\cdot d_{\text{max}}^{2})$
Path Existence Determine whether a specific path exists between two nodes in a given graph. Depth-First Search $O(|V|+|E|)$
Shortest Path Determine the minimum distance between two nodes in a given graph. Dijkstra $O({|V|}^{2}+|E|)$
Table 17: Details of out-toolset Graph Algorithmic Reasoning Tasks.
Task Description Algorithm Time Complexity
Maximum Flow Compute maximum traffic flow from source to sink node. Exclusive for Transportation. Edmonds-Karp $O(|V|\cdot|E|^{2})$
Diameter Calculation Find the longest shortest path in the given graph. Exclusive for Transportation. Floyd-Warshall $O(|V|^{3})$
Max Core Calculation Identify the largest k-core subgraph. Exclusive for Social. Core Decomposition $O(|E|)$
Connected Components Count the number of all connected subgraphs. Exclusive for Social. Breadth-First Search $O(|V|+|E|)$
Common Neighbors Count shared connections between two nodes. Exclusive for Web. Set Intersection $O(\min(d(u),d(v)))$
PageRank Calculation Compute node PageRank based on link structure. Exclusive for Web. Power Iteration $O(k|E|)$
Reference Match Calculate the number of commonly cited papers between two papers. Exclusive for Citation. Set Intersection $O(|A|+|B|)$
Clustering Coefficient Calculate local clustering coefficient of nodes. Exclusive for Citation. Triangle Counting $O(|V|\langle k^{2}\rangle)$
The out-toolset Graph Algorithmic Reasoning Tasks present domain-specific challenges that require specialized algorithmic solutions beyond the toolset. These tasks include Maximum Flow and Diameter Calculation for Transportation, Max Core Calculation and Connected Components for Social, Common Neighbors and PageRank Calculation for Web, and Reference Match and Clustering Coefficient for Citation. Each task employs distinct algorithms with varying computational complexities tailored to their specific domain requirements, evaluating LLMs’ task-specific model generation capabilities.
F.3 Graph Predictive Modeling Tasks
The Graph Predictive Modeling Tasks further evaluate LLMs’ reasoning abilities on graphs by introducing prediction-based challenges that differ from the deterministic results of Structural Querying tasks and Algorithmic Reasoning tasks. These tasks specifically assess LLMs’ ability to leverage neural networks for predictive modeling, with each task explicitly assigned a distinct neural architecture to facilitate standardized evaluation. For Traffic Prediction from PeMS datasets Shuman et al. [2013], we preprocess and store time-series data in Numpy arrays. Similarly, Node Classification tasks from Cora Craven et al. [1998] utilize pre-computed text embeddings stored in PyG format. Unlike the traffic flow prediction task and node classification task which are derived from PeMS and Cora datasets with complete task construction and evaluation schemes, Link Prediction (constructed from SNAP Social Circles McAuley and Leskovec [2012] and Google Web Graph McAuley and Leskovec [2012]) requires first constructing both the graph and the task itself. The Link Prediction tasks employ a specially constructed dataset generated through original graph sampling with added random noise, where the task is specifically defined as evaluating whether two unconnected nodes may have latent relationships based on existing edge patterns. These predictive tasks include Traffic Prediction for Transportation, Link Prediction for both Social and Web domains, and Node Classification for Citation, each implemented with domain-specific neural architectures to assess LLMs’ predictive reasoning capabilities on graph-structured data.
Table 18: Details of Graph Predictive Modeling Tasks.
Task Description Algorithm Time Complexity
Traffic Prediction Predict future traffic flow patterns. Exclusive for Transportation. LSTM $O(T\cdot n^{2})$
Link Prediction Predict missing connections between users. Exclusive for Social. GCN $O(L\cdot|E|\cdot d^{2})$
Link Prediction Predict potential hyperlinks between pages. Exclusive for Web. GraphSAGE $O(\prod_{l=1}^{L}s_{l}\cdot d^{2})$
Node Classification Classify papers into research topics. Exclusive for Citation. GAT $O(|V|\cdot d^{2}+|E|\cdot d)$
Appendix G Limitations
Although GraphCogent outperforms other baselines, it still has several limitations. First, while the framework performs well on static graphs, it does not natively support dynamic graphs with evolving topologies or edge weights. The Buffer Module assumes static data, requiring full reconstruction for updates, which introduces computational overhead. Second, the Model Agent’s reasoning capabilities are currently limited to pre-trained models (e.g., GCN, LSTM) and deterministic tasks. Supporting more advanced models may require integrating more powerful foundation models (such as DeepSeek-V3 or GPT-4o) or additional training. Moreover, the framework struggles with NP-hard problems at larger scales due to computational complexity, and its current performance evaluation primarily relies on accuracy metrics, which may not fully assess tasks requiring both numerical results and path validity (e.g., shortest path or max flow problems). Additionally, while the framework standardizes diverse graph representations, highly specialized formats (e.g., molecular graphs) may require additional domain-specific parsing. These limitations suggest directions for future work, including enhanced support for dynamic graphs, broader model generalization, and more comprehensive evaluation metrics.
Appendix H Error Analysis
In this section, we analyze the errors of GraphCogent (Llama3.1-8B based) within the Graph4real benchmark for both in-toolset tasks and out-toolset tasks.
H.1 In-toolset tasks
For in-toolset tasks, we selected Edge Count, Triangle Count, Cycle Detection, and Path Existence, analyzing error patterns across 500 test instances at the middle scale (100 nodes). The errors were categorized into the following two types:
  • Edge Error: Structural deviations of origin graph and transformed graph caused by different edges during transformation process.
  • Tool Error: Incorrect tool selection or parameterization in the Execution Module
Table 19: Error distribution across in-toolset tasks.
Task Type Edge Error Tool Error Accuracy
Edge Count 19 2 97.8%
Triangle Count 18 1 98.8%
Cycle Detection 21 5 99.0%
Path Existence 19 13 97.4%
The results in Table 19 demonstrate that GraphCogent achieves high accuracy across in-toolset tasks, but performance varies by task type. For edge-dependent tasks (e.g., Edge Count, Triangle Count), errors primarily stem from Edge Errors, structural discrepancies introduced during graph transformation. These tasks rely heavily on precise edge information, making them sensitive to minor deviations in the adjacency list conversion. In contrast, topology-independent tasks (e.g., Cycle Detection, Path Existence) show a reversed pattern. While Edge Errors may occur, their impact is mitigated because task outcomes often do not depend on specific edge subsets. Therefore, Tool Errors dominate the main error types for these tasks.
H.2 Out-toolset tasks
For out-toolset tasks, we select Max Flow on Transportation and Link Prediction on Social, analyzing error patterns across 500 test instances at the middle-scale graph (100 nodes). We classify errors into three categories:
  • Discrimination Error: Occurs when the Reasoning Agent fails to identify the task as out-toolset task, leading to inappropriate tool calls instead of activating the Model Agent.
  • Execution Error: Arises when the task-specific model generated by Model Agent contains logical flaws or implementation issues that prevent successful execution.
  • Format Error: Applies when the task-specific model produces correct results but returns them in non-compliant formats.
Table 20: Error distribution across out-toolset tasks.
Task Type Discrimination Error Execution Error Format Error Accuracy
Max Flow 35 6 11 89.6%
Link Prediction 6 58 10 85.2%
The results are reported in Table 20. For Max Flow, discrimination errors dominate, primarily due to the Reasoning Agent’s failure to discriminate task boundaries when problems exceed the toolset’s coverage. In such cases, the Reasoning Agent incorrectly attempts tool-based solutions despite the task requiring model generation. Format errors are secondary, often manifesting as mismatched outputs (e.g., returning paths instead of the max flow value). In contrast, Link Prediction is primarily affected by execution errors, which stem from internal tensor mismatches in the generated neural models (e.g., dimension inconsistencies in GCN layers or invalid adjacency matrix operations). Both discrimination and format errors occur less frequently, suggesting that the model reliably identifies the task type and produces valid outputs.
Appendix I LLM Prompts
The prompts we used for Sensory Agent are shown in Fig. 12. Prompts for Reasoning Agent are presented in Fig. 10 and Fig. 11. Fig. 13 gives the example of prompt for Model Agent. Fig. 14 gives an example of task generation.
Figure 10: Prompt for Reasoning Agent.
Figure 11: Prompt for Toolsets.
Figure 12: Prompt for Sensory Agent.
Figure 13: Prompt for Model Agent.
Figure 14: Task generation template.
Appendix J GraphCogent Workflow
Algorithm 2 presents the workflow of GraphCogent. For the four key components of the reasoning process, we provide corresponding case examples in Fig. 15, Fig. 16, Fig. 17, and Fig. 18.
Algorithm 2 GraphCogent Framework
1:Graph task $T$, Input graph $G$
2:Task result $R$
3:Sensory Module:
4:Parse raw graph data via subgraph sampling
5:Transform text representations to adjacency lists using heuristic prompts
6:Verify transformation via Graph Verifier (quantity consistency + format compliance)
[…]
8:Buffer Module:
9:Construct complete graph from adjacency lists
10:Transform to multiple representations: NetworkX/PyG/NumPy
11:Build data indices (dimensionality, schema, metadata)
12:Store preprocessed graph data $\tilde{G}$
13:Execution Module:
14:if $T$ is in common toolset coverage then
15:  $\triangleright$ See Fig. 16 for in-toolset reasoning example
16:  Select appropriate tool from common toolset
17:  Retrieve required data format from Buffer Module
18:  Execute tool calling on $\tilde{G}$
19:  Return result $R$
20:else
21:  $\triangleright$ See Fig. 17 for out-toolset reasoning example
22:  Generate task-specific model $M$ for $T$
23:  $\triangleright$ See Fig. 18 for model generation example
24:  Combine $M$ with $\tilde{G}$ from Buffer Module
25:  Execute generated model
26:  Return result $R$
27:end if
Figure 15: Example of Sensory Agent transformation process.
Figure 16: Example of Reasoning Agent’s in-toolset processing.
Figure 17: Example of Reasoning Agent’s out-toolset processing.
Figure 18: Example of Model Agent’s model generation.